# Learning Ordinality in Semantic Segmentation

### v1

@@ 5 Results
To evaluate whether the inclusion of domain knowledge during the training of the models would help the network learn better with scarce data for autonomous driving scenarios, an experiment that consisted of varying the scale of the dataset used to train the models was performed. For each method, the best-performing lambda in the out-of-distribution scenario was selected, the rationale being that those models are the best at generalizing to unseen scenarios, which is useful when training with low amounts of data. For the $L_{\text{CO2}}$ and $L_{\text{CSNP}}$ losses, $\lambda=10.0$ was used, and for the $L_{\text{CSDT}}$, $\lambda=0.1$ was used. Figure 6 shows the Dice coefficient results for the dataset scale variation experiment.
@@
When testing with the BDD10K dataset, the $L_{\text{CO2}}$ and $L_{\text{CSNP}}$ losses achieve better Dice coefficient results at scales $0.25$ and $0.5$ when compared with the cross-entropy baseline, which suggests that, indeed, these losses help the network learn better when data is scarce. Especially when using the CO2 loss, resulting in absolute gains of $1.2\%$ ($5.7\%$ in relative terms) in the Dice coefficient performance at scale $0.25$ over cross-entropy. CO2 continues beating cross-entropy Dice performance through scales $0.1$ and $0.05$.
@@
Cross-entropy generalizes better than the spatial methods when testing with the Cityscapes dataset in an out-of-distribution scenario for lower scales. However, the CO2 loss continues to generalize better than the cross-entropy throughout all scales, achieving a maximum absolute gain of $5.3\%$ ($15.7\%$ in relative terms) in the Dice coefficient at scale $0.75$. Still, it can be seen that the generalization ability of CO2 decreases at a higher rate than its performance on BDD10K – training with fewer data has a higher impact on the model’s generalization ability.
@@
Figure 11: Comparison of the influence of cross-entropy and CO2 losses on model output in out-of-distribution inference. (a) Input image. (b) Ground-truth. (c) Cross-entropy. (d) CO2 ($\lambda=10.0$). The CO2 loss promotes more accurately identification of pedestrians, two-wheel vehicles, and riders when compared with the cross-entropy loss output.
5.1 Discussion
@@
The representation space, i.e., pixel-wise, $L_{\text{CO2}}$ performed well in both the biomedical and autonomous driving datasets, especially when considered in an out-of-distribution domain – it could potentially be applied in a real-world scenario in order to improve the generalization capabilities of perception algorithms.
The structured space, i.e., spatial, $L_{\text{CSNP}}$ and $L_{\text{CSDT}}$ losses, may not apply to the autonomous driving scenario in their current form, at least at high regularization weights. The scene perspective from the car makes it so that there is a large amount of valid contact surface between non-ordinally adjacent classes – in a 2D projection of the real world, the absolute minimization of these contact surfaces may not be the best solution since occlusions and different perspectives may originate legitimate contact between non-ordinally adjacent classes. Relaxed adaptations of these methods that consider this type of contact could be devised in the future. However, in the biomedical datasets, these methods performed well, with their greatest difficulty being the existence of occlusions.
As seen, the choice of $\lambda$ value, i.e., regularization weight, is critical for the performance of the proposed methods. For this reason, in order to be applied to different domains, there should be an empirical study of the influence of the lambda value in the segmentation performance in that specific domain and the choice of a value that is a balance between the ordinal metrics and the Dice coefficient, i.e., a value that promotes some unimodality and spatial consistency but also does not hurt Dice performance to the point where it is unusable.
@@ 6 Conclusion
This paper explored ordinality in the context of structured data, proposing a formal definition of an ordinal model for generic graph structures. The concept was then specifically studied for semantic segmentation, where each image can be interpreted as a graph.
@@
    the segmentation adaptation of the representation $L_{\text{CO2}}$ loss;
@@
In addition, two metrics were proposed to evaluate the network output’s ordinal consistency: (1) the percentage of unimodal pixels and (2) the contact surface between the segmentation masks of non-ordinally adjacent classes.
The proposed methods were initially validated on five biomedical datasets and two autonomous driving datasets, resulting in more ordinally consistent models without significantly impacting the Dice coefficient. To evaluate the methods’ impact on the models’ generalization capability, the resulting models were tested in an autonomous driving out-of-distribution scenario. Furthermore, the autonomous driving models were also tested with scaled-down versions of the BDD100K dataset to evaluate how the network learns with scarce data. The ordinal methods achieved maximum improvements in the Dice coefficient with an absolute value of $5.3\%$ ($15.7\%$ in relative terms) in the out-of-distribution domain.
@@
Future research topics could include: (i) the development of more flexible spatial ordinal segmentation methods, allowing for limited contact between non-ordinally adjacent classes, such as in the case of occlusions and different perspectives; and (ii) the development of novel methods that leverage ordinal constraints not necessarily consisting of augmented loss functions.
Furthermore, the current work could contribute to other types of consistency. The partially ordered domains explored in this work were structural (e.g., a vehicle may be located on the road or in parking spaces), but a representational partially ordered domain could correspond to hierarchical segmentation [22], where each pixel has a taxonomy (e.g., a motorcycle and bicycle are both two-wheels; two-wheels and four-wheels are both types of vehicles).
Also, part segmentation could be considered a different type of structural consistency, where parts of an object are next to each other. Super-pixels have been used to impose constraints, such as the “head” appearing above “upper body” or “hair” being above “head” [23]. Similarly, the function of Restricted Boltzmann Machines has been used to favor certain classes based on their spatial locations [24]. This work could be considered an alternative that promotes such structural consistency using loss penalties.
7 Source Code

### v2

@@ 5 Results
To evaluate whether the inclusion of domain knowledge during the training of the models would help the network learn better with scarce data for autonomous driving scenarios, an experiment that consisted of varying the scale of the dataset used to train the models was performed. For each method, the best-performing lambda in the out-of-distribution scenario was selected, the rationale being that those models are the best at generalizing to unseen scenarios, which is useful when training with low amounts of data. For the $L_{\text{O2}}$ and $L_{\text{CSNP}}$ terms, $\lambda=10$ was used, and for the $L_{\text{CSDT}}$, $\lambda=0.1$ was used. Figure 7 shows the Dice coefficient results for the dataset scale variation experiment.
@@
When testing with the BDD10K dataset, CE+O2 and CE+CSNP regularization terms achieve better Dice coefficient results at scales $0.25$ and $0.5$ when compared with the CE baseline, which suggests that, indeed, these losses help the network learn better when data is scarce. Especially when using the CE+O2 loss, resulting in absolute gains of $1.2\%$ ($5.7\%$ in relative terms) in the Dice coefficient performance at scale $0.25$ over CE. CE+O2 continues beating CE Dice performance through scales $0.1$ and $0.05$.
@@
Cross-entropy alone generalizes better than when used in conjunction with the ordinal regularizers when testing with the Cityscapes dataset in an out-of-distribution scenario for lower scales. However, CE+O2 continues to generalize better than the cross-entropy throughout all scales, achieving a maximum absolute gain of $5.3\%$ ($15.7\%$ in relative terms) in the Dice coefficient at scale $0.75$. Still, it can be seen that the generalization ability of CE+O2 decreases at a higher rate than its performance on BDD10K – training with fewer data has a higher impact on the model’s generalization ability.
@@ 5.1 Discussion
Figure 12: Comparison of the influence of cross-entropy and O2 term on model output in out-of-distribution inference. (a) Input image. (b) Ground-truth. (c) CE alone. (d) CE+O2 ($\lambda=10.0$). The O2 term promotes more accurate identification of pedestrians, two-wheel vehicles, and riders when compared with the cross-entropy loss output.
@@
The representation space, i.e., pixel-wise, $L_{\text{O2}}$ performed well in both the biomedical and autonomous driving datasets, especially when considered in an out-of-distribution domain – it could potentially be applied in a real-world scenario in order to improve the generalization capabilities of perception algorithms.
The structured space regularization terms, $L_{\text{CSNP}}$ and $L_{\text{CSDT}}$, may not apply to the autonomous driving scenario in their current form, at least at high regularization weights. The scene perspective from the car makes it so that there is a large amount of valid contact surface between non-ordinally adjacent classes – in a 2D projection of the real world, the absolute minimization of these contact surfaces may not be the best solution since occlusions and different perspectives may originate legitimate contact between non-ordinally adjacent classes. Relaxed adaptations of these methods that consider this type of contact could be devised in the future. However, in the biomedical datasets, these methods performed well, with their greatest difficulty being the existence of occlusions.
As seen, the choice of $\lambda$ value, i.e., regularization weight, is critical for the performance of the proposed methods. For this reason, in order to be applied to different domains, there should be an empirical study of the influence of the lambda value in the segmentation performance in that specific domain and the choice of a value that is a balance between the ordinal metrics and the Dice coefficient, i.e., a value that promotes some unimodality and structural consistency but also does not hurt Dice performance to the point where it is unusable.
@@ 6 Conclusion
This paper addresses the overlooked role of ordinality in semantic segmentation by introducing methods that enforce both pixel-level and structural ordinal consistency. By leveraging the inherent relationships between classes, we developed two regularization loss functions and an evaluation metric tailored to promote ordinal consistency in structured image data.
@@
    a segmentation adaptation of the representation $L_{\text{O2}}$ loss which penalizes ordinal inconsistencies within each pixel;
@@
In addition, two metrics were proposed to evaluate the network output’s ordinal consistency: (1) the percentage of unimodal pixels and (2) the contact surface between the segmentation masks of non-ordinally adjacent classes. The first evaluates in-pixel consistency, while the second evaluates spatial consistency.
The proposed methods were initially validated on five biomedical datasets and two autonomous driving datasets, resulting in more ordinally consistent models without significantly impacting the Dice coefficient. The resulting models were tested in an autonomous driving out-of-distribution scenario to evaluate the methods’ impact on the models’ generalization capability. Furthermore, the autonomous driving models were also tested with scaled-down versions of the BDD100K dataset to evaluate how the network learns with scarce data. The ordinal methods achieved maximum improvements in the Dice coefficient with an absolute value of $5.3\%$ ($15.7\%$ in relative terms) in the out-of-distribution domain.
@@
Future research topics could include: (i) the development of more flexible structural ordinal segmentation methods, allowing for limited contact between non-ordinally adjacent classes, such as in the case of occlusions and different perspectives; and (ii) the development of novel methods that leverage ordinal constraints not necessarily consisting of augmented loss functions.
Furthermore, the current work could contribute to other types of consistency. The partially ordered domains explored in this work were structural (e.g., a vehicle may be located on the road or in parking spaces), but a representational partially ordered domain could correspond to hierarchical segmentation [11], where each pixel has a taxonomy (e.g., a motorcycle and bicycle are both two-wheels; two-wheels and four-wheels are both types of vehicles). Also, part segmentation [12, 13] could be considered a different type of structural consistency, where parts of an object are next to each other (e.g., the “head” appearing above “upper body” or “hair” above “head”).
Appendix A Source Code
The full source code is provided in https://github.com/rafaavc/ordinal-semantic-segmentation. The code for the proposed spatial losses is small and is presented here in full: